OpenStudy (anonymous):

x^2+y^2=9 16x^2-4y^2=64

OpenStudy (anonymous):

are you supposed to solve for x and y?

OpenStudy (anonymous):

Solving for x^2 and y^2, using linear algebra: u = x^2, v = y^2 (i) u + v = 9 (ii) 3u - 2v = 32 [dividing your second equation by 2] (iii) 2u + 2v = 18 [doubling equation (i)] 5u = 50 [adding equations (ii) and (iii)] u = 10 10 + v = 9 v = -1. Then x = +- sqrt(10), y = +-i. If you were to graph the circle and the hyperbola in the same x-y plane, you would see that they do not intersect anywhere.